OpenStudy (anonymous):

Find all solutions in the interval [0, 2π). 2 sin^2x = sin x

OpenStudy (anonymous):

double angle formula set equal to zero factor solve for x

OpenStudy (anonymous):

i dont know how to use the double angle formula @completeidiot

OpenStudy (anonymous):

its a trig identity \[\sin(2x) = 2sinxcosx\]

OpenStudy (anonymous):

ok and what do i do after that?

OpenStudy (anonymous):

substitute.... then follow above steps

OpenStudy (anonymous):

substitute what ? @completeidiot

OpenStudy (anonymous):

double angle formula...

OpenStudy (anonymous):

could you explain a little more on how to do that ? @completeidiot

OpenStudy (anonymous):

oh crap, misread the problem, you dont need to use double angle formula

OpenStudy (anonymous):

just set equal to zero, factor, then solve

OpenStudy (anonymous):

If your problem was 2sin(2x) = sin x then substituting using double angle formula would be 2 ( 2sin x cos x)) = sin x

OpenStudy (anonymous):

ok so \[2SIN^2X-SINX=0\] ?

OpenStudy (anonymous):

now factor

OpenStudy (anonymous):

factor what ? @completeidiot

OpenStudy (anonymous):

look at the equation and tell me if you are capable of factoring anything out

OpenStudy (anonymous):

from my equation i got \[2\sin^2x-sinx=0\] and then i got \[sinx(2\sin-1)\] @completeidiot

OpenStudy (anonymous):

right and that expression is equal to zero now use the zero property of multiplication basically it means either sinx =0 or 2 sin x -1 = 0 solve for x in both cases

OpenStudy (anonymous):

i got x=0 and sinx=1/2 @completeidiot

OpenStudy (anonymous):

just saying, there should be 4 different answers i suggest looking a unit circle

OpenStudy (anonymous):

so im guessing the answers are 0, pie, pie/6 , & 5pie/6 ? @completeidiot

OpenStudy (anonymous):

looks about right